# Linear Combination of Saved Checkpoints Makes Consistency and Diffusion Models Better

### v2

@@ 0.C.1.3 The Critical Role of Search Space Configuration
0.C.2 Discussion on the Training Variance of DM and CM
As discussed in Sec. 3, we are motivated by the observation that the training objectives of DM and CM tend to introduce substantial variance in gradient estimations. We further discuss the potential reasons or hypothesis of such phenomenon as follows.
0.C.2.1 Unbiased Estimation as Objective
The objectives used in DM and CT are not accurate for any single batch. Instead, they serve as an unbiased estimation. Specifically, for DM, the neural network learns the score function at any $(\bm{x}_{t},t)$, denoted as $\bm{s}_{\theta}(\bm{x}_{t},t)$. The ground truth score function $\nabla\log p_{t}(\bm{x}_{t})$ is given as [40]:
@@ 0.C.2.1 Unbiased Estimation as Objective
This formula implies that every sample $\bm{y}$ in the dataset contributes to the ground truth score function at any $(x_{t},t)$. However, the training objective Eq. 3 provides only an unbiased estimation of the ground truth score Eq. 9. At every iteration, the current target for the network is not identical to the ground truth score function but is determined by the sample $\bm{y}$ randomly drawn from the dataset and the noise level $\sigma_{t}$. Thus, even if $\bm{s}_{\theta}(\bm{x}_{t},t)$ exactly matches $\nabla p_{t}(\bm{x}_{t})$ for $\forall(\bm{x}_{t},t)$, its gradient estimation using a mini-batch of data does not converge to zero. The approximation of ground truth score function can only be obtained through the expectation over many training iterations. For CT, the situation is very similar. To simulate the current ODE solution step without a teacher diffusion model, Song et al. [39] utilize the Monte Carol estimation $-\frac{\bm{x}_{t}-\bm{y}}{t^{2}}$ to replace the ground truth score function $\nabla\log p_{t}(\bm{x}_{t})=\mathbb{E}_{\bm{y}\sim p_{\text{data}},\bm{x}_{t}\sim\mathcal{N}(\bm{y},t^{2}\mathbf{I})}(-\frac{\bm{x}_{t}-\bm{y}}{t^{2}}|\bm{x}_{t})$ under the noise schedule from EDM [15] as denoted in Eq. 5. Therefore, the objective is also not accurate for any single batch of data and the model has to learn to fit the target model output after solving the current step with the expectation of score function among many training iterations. These types of objectives, which are not accurate for any single batch, introduce high variance to the gradient estimation.
0.C.2.2 Model across Different Time steps
For both DM and CM, the neural network has to be trained at various time steps. Since the model input and output follow different distributions, the gradients at different time steps may conflict with each other. This phenomenon of negative transfer between timesteps has been studied in several previous work, where they train different diffusion neural networks at different timesteps and achieve better performance [6, 5]. The misalignment between the objectives of different time steps may also contribute to the high variance in the gradient estimation.
0.C.2.3 Error Accumulation of CM
For CM, the model learns to approximate the output of the target model at the previous step. This could potentially introduce the problem of error accumulation [1]. Consequently, any noise introduced during training at early time steps is likely to lead to inaccuracies in the target model, which may be magnified in subsequent timesteps. This property of CM amplifies the high training variance, particularly for one-step sampling.
Appendix 0.D Future Work
@@
Appendix 0.E Visualization
In this section, visualizations of images generated by DM (Figs. 9 and 10), CD (Figs. 11 and 12), and CT (Figs. 13 and 14) are presented. Many images produced by the EMA model are observed to be similar to those produced by the model derived from our LCSC when using the same noise input. This is expected given that both models are based on weights from the same training cycle. To more clearly highlight the distinctions between EMA and LCSC, we generate 50K images from EMA and LCSC using the same set of noise inputs, order the image pairs according to their Euclidean distances in the inception feature space, and randomly select the images with large distances. In general, images generated by the LCSC model are found to display enhanced sharpness, diminished noise, and more distinct object representation.

### v3

@@ F.1.3 The Critical Role of Search Space Configuration
Figure 13: Visualization of coefficient pattern at different number of search iterations. The experiment is conducted on ImageNet using CT.
F.1.4 Formation of the Search Pattern
We observe that the search pattern is established during the early stage of the search, while the later stage primarily amplifies it to a certain magnitude, as shown in Fig. 13.
F.2 Regularized Evolutionary Search
We conduct additional experiments to evaluate the impact of regularization. Specifically, during the search process, we clip all coefficients to be below 1, thereby constraining the search to a restricted space. Using the same initialization and random seed, Fig. 14 shows that LCSC produces smaller and more homogeneous coefficients with regularization compared to those obtained without it. However, as noted in the figure caption, applying regularization results in worse performance, suggesting that constraining LCSC to smaller and more homogeneous coefficients is not advantageous.
(a) Visualization of searched coefficients with regularization.
(b) Visualization of searched coefficients without regularization.
Figure 14: Visualization of searched coefficients with and without regularization. Respective FID/IS/Prec/Rec are: 13.40/33.4/0.64/0.55 (left), 10.5/36.8/0.66/0.56 (right).
F.3 Convergence Curve of LCSC
Fig. 15(a) illustrates the convergence curve of the models searched by LCSC. At each point, corresponding to a specific number of training iterations, we perform LCSC with 2K search iterations using the most recent checkpoints available at that point. The results indicate that as the number of training iterations increases, the models searched by LCSC converge to progressively lower FID values. Notably, the convergence curve of LCSC exhibits a similar trend to that of the EMA model’s convergence curve. However, LCSC consistently achieves lower FID values compared to the EMA model, highlighting its enhanced ability to effectively merge historical checkpoints. Additionally, Fig. 15(b) shows the convergence curve of LCSC iteself. This curve consistently reduces as the number of search iteration increases.
(a) Training convergence curve with LCSC.
(b) Search convergence curve of LCSC.
Figure 15: Convergence curves of LCSC on CIFAR-10 with CD.
F.4 Discussion on the Training Variance of DM and CM
In Sec. 3, we mentioned that the training objectives of DM and CM tend to introduce substantial variance in gradient estimations. We further discuss the potential reasons or hypothesis of such phenomenon as follows.
F.4.1 Unbiased Estimation as Objective
The objectives used in DM and CT are not accurate for any single batch. Instead, they serve as an unbiased estimation. Specifically, for DM, the neural network learns the score function at any $(\bm{x}_{t},t)$, denoted as $\bm{s}_{\theta}(\bm{x}_{t},t)$. The ground truth score function $\nabla\log p_{t}(\bm{x}_{t})$ is given as (Song et al., 2020b):
@@ F.4.1 Unbiased Estimation as Objective
This formula implies that every sample $\bm{y}$ in the dataset contributes to the ground truth score function at any $(x_{t},t)$. However, the training objective Eq. 2 provides only an unbiased estimation of the ground truth score Eq. 78. At every iteration, the current target for the network is not identical to the ground truth score function but is determined by the sample $\bm{y}$ randomly drawn from the dataset and the noise level $\sigma_{t}$. Thus, even if $\bm{s}_{\theta}(\bm{x}_{t},t)$ exactly matches $\nabla p_{t}(\bm{x}_{t})$ for $\forall(\bm{x}_{t},t)$, its gradient estimation using a mini-batch of data does not converge to zero. The approximation of ground truth score function can only be obtained through the expectation over many training iterations. For CT, the situation is very similar. To simulate the current ODE solution step without a teacher diffusion model, Song et al. (Song et al., 2023) utilize the Monte Carol estimation $-\frac{\bm{x}_{t}-\bm{y}}{t^{2}}$ to replace the ground truth score function $\nabla\log p_{t}(\bm{x}_{t})=\mathbb{E}_{\bm{y}\sim p_{\text{data}},\bm{x}_{t}\sim\mathcal{N}(\bm{y},t^{2}\mathbf{I})}(-\frac{\bm{x}_{t}-\bm{y}}{t^{2}}|\bm{x}_{t})$ under the noise schedule from EDM (Karras et al., 2022) as denoted in Eq. 4. Therefore, the objective is also not accurate for any single batch of data and the model has to learn to fit the target model output after solving the current step with the expectation of score function among many training iterations. These types of objectives, which are not accurate for any single batch, introduce high variance to the gradient estimation.
F.4.2 Model across Different Time steps
For both DM and CM, the neural network has to be trained at various time steps. Since the model input and output follow different distributions, the gradients at different time steps may conflict with each other. This phenomenon of negative transfer between timesteps has been studied in several previous work, where they train different diffusion neural networks at different timesteps and achieve better performance (Hang et al., 2023; Go et al., 2024). The misalignment between the objectives of different time steps may also contribute to the high variance in the gradient estimation.
F.4.3 Error Accumulation of CM
For CM, the model learns to approximate the output of the target model at the previous step. This could potentially introduce the problem of error accumulation (Berthelot et al., 2023). Consequently, any noise introduced during training at early time steps is likely to lead to inaccuracies in the target model, which may be magnified in subsequent timesteps. This property of CM amplifies the high training variance, particularly for one-step sampling.
Appendix G Future Work
@@
Appendix H Visualization
In this section, visualizations of images generated by LCM-LoRA (Fig. 16), DM (Figs. 17 and 18), CD (Figs. 19 and 20), and CT (Figs. 21 and 22) are presented. Many images produced by the EMA model are observed to be similar to those produced by the model derived from our LCSC when using the same noise input. This is expected given that both models are based on weights from the same training cycle. To more clearly highlight the distinctions between EMA and LCSC, we generate 50K images from EMA and LCSC using the same set of noise inputs, order the image pairs according to their Euclidean distances in the inception feature space, and randomly select the images with large distances. In general, images generated by the LCSC model are found to display enhanced sharpness, diminished noise, and more distinct object representation as well as details.
(a) Vanilla training
(b) LCSC
(c) Vanilla training
(d) LCSC
(e) Vanilla training
(f) LCSC
(g) Vanilla training
(h) LCSC
Figure 16: Examples of the images generated by LCSC and vanilla training. The prompts for the images are ’A large passenger airplane flying through the air‘, ’A brown purse is sitting on a green bench‘, ’A man getting a drink from a water fountain that is a toilet‘, ’A picture of a man playing a violin in a kitchen‘ respectively.